OpenStudy (anonymous):

What is the surface area of a cube with side length 9 kilometers? two hundred sixteen square kilometers three hundred twenty-four square kilometers four hundred eighty-six square kilometers seven hundred twenty-nine square kilometers

OpenStudy (anonymous):

k if one side is something then everything is that something because its a cube correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how many side lengths does a cube have?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

i mean 2

OpenStudy (anonymous):

[]

OpenStudy (anonymous):

k no and no

OpenStudy (anonymous):

we know its a cube right?

OpenStudy (anonymous):

so in this case you can simply cube your number

OpenStudy (wolf1728):

Each face = 9 * 9 = 81 kilometers² All six faces = 6 * 81 = 486 kilometers²